OpenStudy (nightmarenight):

Solve the equation. x2 = 160

OpenStudy (nightmarenight):

possible answers:

OpenStudy (mathstudent55):

\(x^2 = 160\) An equation of the form \(X^2 = k\), where X is an expression in x, and k is a number can be solved by writing it as \(X = -\sqrt{k}\) or \( X = \sqrt{k} \) and solving each equation.